OpenStudy (anonymous):

Alight I got my answers but dont know which to choose . dentify the correct x intercept and y intercept of the equation x - 2y = 8.

OpenStudy (anonymous):

I got 4 and 8 x-0=8 x=8 <--- x-2y=8 0-2y=8 y=4 <---

OpenStudy (anonymous):

x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (mayaal):

Statement 2 is correct :x intercept = (8, 0); y intercept = (0, -4)

OpenStudy (anonymous):

thnkz #300 I picked that one but i wasent sure if was gonna be right .

OpenStudy (mayaal):

U r Welcome....!